Cost of producing Product X Cost of producing


Country A $10 $8


Country B $20 $5


(i) Without a trade what is the cost of producing product X for


country A and which one for B? (15 Marks)


(ii) Which country has a comparative advantage for product X and


product Z and why? (15 Marks)


(iii) Which country has an absolute advantage for both products and


why? (15 Marks)


(iv) If the two countries trade with each other which product will


prefer to export and why? (15 Marks)

1
Expert's answer
2022-02-01T10:16:13-0500

X Z

Country A $10 $8

Country B $20 $5

i)Without a trade, the cost of producing X in country A : the cost producing X in country B

i.e 10:20

"\frac{10}{20}=" "\frac{1}{2}=0.5"

ii)for country A,"(\frac{a_x}{b_x})_A=\frac{10}{20}=\frac{1}{2}=0.5"

for country B, ("\frac{a_z}{b_z})_B=\frac{8}{5}=1.6"

Country A has a comparative advantage due to the low opportunity cost.

iii) Country B has an absolute advantage due to the high opportunity cost.

iv)iv)The H-O theorem states that a country exports those goods that makes intensive use of factors that are locally abundant hence if the two countries are to trade, the product X will be exported

because it has a low opportunity cost and is a capital intensive good,